# New Job, New Gender? Measuring the Social Bias in Image Generation Models

### v1

@@ 4.4. RQ3 - Bias Mitigation
Specifically, we select the top biased profession words shown in Table 3, add an additional system prompt that ”maintains the same gender/race/age as the input image” and then regenerate the images. Then, we compare the bias score when generating with the original prompt (denoted as ”Ori”) to the bias score when generating with the additional system prompt (denoted as ”Miti”). As the results are shown in Table 5, the average bias score when generating with the additional prompt is relatively smaller (e.g. 0.40 v.s. 0.98 for SD1.5), indicating that adding this specific prompt can reduce the bias to a certain extent, but is far from completely eliminating it.
Table 5. Bias Mitigation Results on Profession
@@
Word Ori Miti Word Ori Miti Word Ori Miti Ori Miti
SD1.5 Secretary 1.00 0.94 Artist 1.24 -0.20 Electrician 1.42 -0.06 0.98 0.40
Taxi Driver -0.67 0.44 Model -0.89 -0.77 Astronomer -0.67 0.01
SD2.1 Nurse 1.00 1.00 Artist 1.23 1.35 Estate Agent 0.66 0.94 0.90 0.79
Soldier —1.0 0.67 Receptionist -0.33 0.69 Doctor -1.17 0.12
SDXL Nurse 0.89 0.39 Economist 1.01 0.22 Receptionist 1.34 1.01 0.98 0.44
Electrician —1.0 0.33 Hairdresser -1.15 -0.63 Fisherman -0.50 -0.06
Midj Nurse 0.78 0.67 Farmer 0.66 0.65 Estate Agent 1.01 -0.26 0.86 0.65
Pilot —1.0 1.0 Lawyer -0.51 0.41 Taxi Driver 1.2 -0.88
P2P Estate Agent 1.00 0.67 Farmer 0.4 -0.02 designer 1.43 0.33 1.08 0.85
Fisherman —0.78 0.67 Plumber -0.70 —0.70 Bartender -2.18 -2.71
Answer to RQ3: BiasPainter can provide insight into bias mitigation methods and also be adopted as an automatic evaluation method for measuring the effectiveness of bias mitigation methods.
5. Threats to Validity
The validity of this work may be subject to some threats. The first threat lies in the AI techniques adopted by BiasPainter for bias identification. Due to the imperfect nature of AI techniques, the biases identified by BiasPainter may be false positives, or BiasPainter may miss some biased generation, leading to false negatives. To relieve this threat, BiasPainter calls commercial-level APIs and deploys complicated pipelines to analyze the race, gender, and age properties, aiming to ensure the soundness. In addition, we also conducted human annotation to show that BiasPainter can achieve high accuracy (i.e., 90.8%) in detecting bias.
The second threat is that the input data of BiasPainter are predefined, both seed images and prompt lists, which may hurt the comprehensiveness of the testing results. To mitigate this threat, we collected diverse and comprehensive seed images and prompt words, all of which are from the real world on the Internet and manually annotated by researchers. Also, we want to highlight that what we provide is a workflow: select seed images, design prompt lists, generate test cases, and find social bias. If users are willing to evaluate other properties on other images, they could follow this workflow to add more images and prompts.
The third threat lies in the image generation models under test in the evaluation. We do not evaluate the performance of BiasPainter on other systems. To mitigate this threat, we chose to test the most widely used commercial image generation software and SOTA academic models provided by famous organizations. In the future, we could test more commercial software and research models to further mitigate this threat.
@@ 6. Related Work
6.1. Testing of AI Software
AI software has been adopted in various domains, such as autonomous driving and face recognition. However, AI software is not robust enough and can generate erroneous outputs that lead to fatal accidents (Ziegler, 2016; Levin, 2018). To this end, researchers have proposed a variety of methods to generate adversarial examples or test cases to measure the reliability of AI software (Carlini et al., 2016; Tu et al., 2021; Luo et al., 2021; Pei et al., 2017; Zhang et al., 2022; Riccio et al., 2020; Humbatova et al., 2021; Pham et al., 2021; Wang et al., 2021; Ma et al., 2018; Tian et al., 2017; Zhang et al., 2019; Zhou et al., 2020; Wang et al., 2023b, a).
NLP software, such as machine translation software and chatbot, has also been widely used in human life. Similar to AI software, researchers have proposed variance methods to validate the reliability of NLP software on the correctness (Gupta, 2020; He et al., 2021; tse Huang et al., 2022; Shen et al., 2022), toxicity (Wang et al., 2023e, f), fairness (Zhang and Harman, 2021; Wan et al., 2023; Wang et al., 2023c).
In contrast to the aforementioned AI software, image generation models serve as a new multimodal AI software that emerged recently, posing challenges for researchers to precisely and comprehensively evaluate and test.
6.2. Bias and Fairness in AI Software
Bias and fairness have gained significant attention in the SE community from various perspectives, such as bias testing (Mehrabi et al., 2019) and bias mitigation (Hort et al., 2022), and for various kinds of software, such as nature language processing models (Sun et al., 2019), recommendation systems (Pitoura et al., 2021), chatbot (Wan et al., 2023), and autonomous driving car (Li et al., 2023).
As one of the most popular AI software, the image generation model is widely used with a sufficient amount of active users. However, to the best of our knowledge, the comprehensive testing framework for the bias and fairness in image generation models remains under investigation. In this paper, we focus on image generation models and propose the first automatic and comprehensive testing framework.
6.3. Bias in Image Generation Models
We systematically reviewed papers on testing the biases in image generation models across related research areas, including software engineering, computer vision, natural language processing, and security.
(Bianchi et al., 2022) is an early work that conducts an empirical study to show the stereotypes learned by text-to-image models. They design different prompts as input and use human annotators to find biased images, without proposing an automatic framework that can trigger the social bias. Inspired by this, (Cho et al., 2022) proposed an automatic framework to evaluate the bias in image generation models. However, their automatic evaluation method failed to accurately detect the bias according to their human evaluation. Also, the generated images are highly biased toward white people so their framework cannot analyze the bias in other groups. More recently, (Wang et al., 2023d) study the gender stereotype in occupations, but the scope of which is limited.
@@ 7. Conclusion
In this paper, we design and implement BiasPainter, a metamorphic testing framework for measuring the social biases in image generation models. Unlike existing frameworks, which only use sentence descriptions as input and evaluate the properties of the generated images, BiasPainter adopts an image editing manner that inputs both seed images and sentence descriptions to let image generation models edit the seed image and then compare the generated image and seed image to measure the bias. We conduct experiments on five widely deployed commercial software and famous research models to verify the effectiveness of BiasPainter. and demonstrate that BiasPainter can effectively trigger a massive amount of biased behavior with high accuracy. In addition, we demonstrate that BiasPainter can help mitigate the bias in image generation models.
Data AVAILABILITY
All the code, data, and results have been uploaded [^17] https://drive.google.com/drive/folders/1VDe5EKszv9TEvJRygK7tIyQDLaeE4Rsn?usp=drive_link and will be released for reproduction and future research.

### v2

@@ 4.4. RQ3 - Bias Mitigation
Specifically, we select the top biased profession words in Table 2, add an additional system prompt that ”maintains the same gender/race/age as the input image” and then regenerate the images. Then, we compare the bias score when generating with the original prompt (denoted as ”Ori”) to the bias score when generating with the additional system prompt (denoted as ”Miti”). As the results are shown in Table 4, the average bias score when generating with the additional prompt is relatively smaller (e.g. 0.40 v.s. 0.98 for SD1.5), indicating that adding this specific prompt can reduce the bias to a certain extent, but is far from completely eliminating it.
Table 4. Bias Mitigation Results Evaluated by BiasPainter
@@
Ori Miti Ori Miti Ori Miti Ori Miti
SD1.5 1.00 0.94 1.24 -0.20 1.42 -0.06 0.98 0.40
-0.67 0.44 -0.89 -0.77 -0.67 0.01
SD2.1 1.00 1.00 1.23 1.35 0.66 0.94 0.90 0.79
—1.0 0.67 -0.33 0.69 -1.17 0.12
SDXL 0.89 0.39 1.01 0.22 1.34 1.01 0.98 0.44
—1.0 0.33 -1.15 -0.63 -0.50 -0.06
Midj 0.78 0.67 0.66 0.65 1.01 -0.26 0.86 0.65
—1.0 1.0 -0.51 0.41 1.2 -0.88
Dalle2 0.44 0.22 0.36 0.06 - - 0.73 0.32
—0.44 -0.44 -0.78 0.14 1.62 0.75
P2P 1.00 0.67 0.4 -0.02 1.43 0.33 1.08 0.85
—0.78 0.67 -0.70 —0.70 -2.18 -2.71
@@ 5. Related Work
Bias and fairness have gained significant attention in the AI community from various perspectives, such as bias measurment (Mehrabi et al., 2019; Garimella et al., 2021; Wang et al., 2023a) and bias mitigation (Hort et al., 2022; Sun et al., 2019a), and for various kinds of AI models, such as nature language processing models (Sun et al., 2019b; Liang et al., 2021; Sap et al., 2019), recommendation systems (Pitoura et al., 2021; Liu et al., 2022), chatbot (Wan et al., 2023), and vision-language pertaining models (Zhang et al., 2022; Ross et al., 2020; Ruggeri et al., 2023).
As one of the most popular AI models, the image generation model is widely used with a sufficient amount of active users. We systematically reviewed papers on evaluating the biases in image generation models across related research areas. (Bianchi et al., 2022) is an early work that conducts an empirical study to show the stereotypes learned by text-to-image models. They design different prompts as input and use human annotators to find biased images, without proposing an automatic framework that can trigger the social bias. Inspired by this, (Cho et al., 2022) proposed an automatic framework to evaluate the bias in image generation models. However, their automatic evaluation method failed to accurately detect the bias according to their human evaluation. Also, the generated images are highly biased toward white people so their framework cannot analyze the bias in other groups. More recently, (Wang et al., 2023b) study the gender stereotype in occupations, but the scope of which is limited.
@@ 6. Conclusion
In this paper, we design and implement BiasPainter, an evaluation framework for measuring the social biases in image generation models. Unlike existing frameworks, which only use sentence descriptions as input and evaluate the properties of the generated images, BiasPainter adopts an image editing manner that inputs both seed images and sentence descriptions to let image generation models edit the seed image and then compare the generated image and seed image to measure the bias. We conduct experiments on six famous image generation models to verify the effectiveness of BiasPainter. and demonstrate that BiasPainter can effectively trigger a massive amount of biased behavior with high accuracy. In addition, we demonstrate that BiasPainter can help mitigate the bias in image generation models.
Acknowledgement
The work described in this paper was supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14206921 of the General Research Fund).